In order to join a voice lesson class you pay 1,500.00 pesos fee then 500 for each class you to go what is the average cost per class? grapgh the function to show whether it forms a straight line or a curve.
1
Expert's answer
2020-10-19T15:50:50-0400

pesos fee = 1500

each class fee = 500

no of class = "x"

average fees(A) per class = "\frac{1500+500x}{x}"

if "x" = 1

average fees = "\frac{1500+500(1)}{1}" = 2000

if "x" = 2

average fees = "\frac{1500+500(2)}{2}" = 1250

if "x" = 3

average fees = "\frac{1500+500(3)}{3}" = 1000

if "x" = 4

average fees ="\frac{1500+500(4)}{4}" = 875

now we got number of classes and average fess , now we plot the graph between these points

"x" = number of classes and A = average fees

"x" = 1 , A = 2000

"x" = 2 , A = 1250

"x" = 3 , A = 1000

"x" = 4 , A = 875